the measure of the second angle of a triangle is seven more than twice the measure of the first. the third angle is fifteen less than the first. find the meausres of all of the angles.
Answer by mananth(16949)

let the first angle be x deg
The second angle = 2x+7
Third angle = x-15
The sum of all 3 angles = 180 deg.
x+2x+7+x-15= 180
4x-8 =180
4x=188
x= 47 deg
2x+7= 101 deg
Third angle = 32 deg
Ananth